OpenStudy (anonymous):

Which system has (1, 3) as the solution? Which system has infinite solutions? Which system has (-8, 0) as the solution? Which system has no solutions? Which system has (-2, -2) as the solution?

OpenStudy (anonymous):

A) x + y = -1 4x + 4y = -4

OpenStudy (anonymous):

B) x - 4y = -8 2x - 3y = -16

OpenStudy (anonymous):

C) x + y = -4 -3x + 2y = 2 D) 4x + 6y = -12 2x + 3y = 9 E) x - y = -2 3x - y = 0

OpenStudy (anonymous):

A) has infinite solutions

OpenStudy (anonymous):

B) has (-8,0) as a solution

OpenStudy (anonymous):

C) has (-2,-2) as the solution

OpenStudy (anonymous):

D) has no solutions

OpenStudy (anonymous):

e has Which system has (1, 3) as the solution?

OpenStudy (anonymous):

E) has (1,3) as a solution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks